For which values of x do y^2=5x^2-4x and y=2x-3 intersect?
Answer by lwsshak3(11628)

For which values of x do y^2=5x^2-4x and y=2x-3 intersect?
**
y^2=5x^2-4y
y=2x-3
..
substitute (2x-3) for y in first equation
(2x-3)^2=5x^2-4x
4x^2-12x+9=5x^2-4x
x^2+8x-9=0
(x+9)(x-1)=0
values of x at intersection: -9 and 1
..
for x=-9
y=2x-3=-18-3=-21
..
for x=1
y=2x-3=2-3=-1
..
2 points of intersection:
(-9,-21) and (1,-1)